### v2

# Dynamical stability of evolutionarily stable strategy in asymmetric games

Abstract
Evolutionarily stable strategy (ESS) is the defining concept of evolutionary game theory. It has a fairly unanimously accepted definition for the case of symmetric games which are played in a homogeneous population where all individuals are in same role. However, in asymmetric games, which are played in a population with multiple subpopulations (each of which has individuals in one particular role), situation is not as clear. Various generalizations of ESS defined for such cases differ in how they correspond to fixed points of replicator equation which models evolutionary dynamics of frequencies of strategies in the population. Moreover, some of the definitions may even be equivalent, and hence, redundant in the scheme of things. Along with reporting some new results, this paper is partially indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games. We present the definitions coherently and scrutinize them closely while establishing equivalences—some of them hitherto unreported—between them wherever possible. Since it is desirable that a definition of ESS should correspond to asymptotically stable fixed points of replicator dynamics, we bring forward the connections between various definitions and their dynamical stabilities. Furthermore, we find the use of principle of relative entropy to gain information-theoretic insights into the concept of ESS in asymmetric games, thereby establishing a three-fold connection between game theory, dynamical system theory, and information theory in this context. We discuss our conclusions also in the backdrop of asymmetric hypermatrix games where more than two individuals interact simultaneously in the course of getting payoffs.
Evolutionarily stable strategy, asymmetric games, replicator equation, dynamical stability, hypermatrix games, relative entropy
I Introduction
Evolutionarily stable strategy (ESS) is the most fundamental concept of evolutionary game theory Smith and Price (1973); Smith (1982); Thomas (1985), and its exploration has yielded a substantial body of scientific literature Smith (1982); Nowak (2006); Hofbauer and Sigmund (1998); Cressman (2003); Weibull (1997). In fact, one may remark Dawkins (1989) that this concept is only second to the Darwin’s contribution in advancing the evolutionary theory. No wonder that the paper Maynard Smith (1974) putting forward this concept has been included as one of eighty-eight foundational papers in complexity science Krakauer (2024) of the 20th century. In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternate strategy (called mutant strategy) cannot invade the population. Essentially, the concept of ESS can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
The ESS was originally introduced for a population where all the individuals are identical (or have same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals. However, the concept of ESS in asymmetric games can be quite tricky. Asymmetric games refer to games that involve individuals categorized into more than one class—each with a distinct role—and the strategic interaction involves interaction between players in different roles. To complicate the matter further, in these games, players belonging to different roles may also have different strategy sets. If there is any interaction among players of same role, the interaction may be termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific. In this context, in this paper we follow the terminology that an asymmetric game may be called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present. The purpose of this paper is to critically discuss various extant definitions of ESS in bimatrix and asymmetric games. Note that henceforth, the phrase ‘asymmetric game’ will not include ‘bimatrix game’ which will be used as a separate phrase. Moreover, unless otherwise specified, we shall work assuming only two roles present in a population for simplicity and without much loss of generality.
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of 2ESS, again with a view to allow for emergence of mixed ESS. The concept is termed Nash–Pareto pair.
The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points. It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases. Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
The outline of the paper is as follows: We commence with Sec. II by outlining the existing definitions of ESS in bimatrix games. Rather than following a chronological description, we adopt a logical and coherent narrative suited to the goal of this paper. We demonstrate the equivalences between some of the definitions. Next, in Sec. III, we highlight caveats and issues associated with these definitions in bimatrix games. We scrutinize the concept of ESS further from the angle of dynamical stability under replicator dynamics. Sec. IV presents the implications on bimatrix games due to intra-specific interactions which essentially turns them into general asymmetric games. In Sec. V, we discuss some interesting information-theoretic connections of ESS. In the next section, Sec. VI, we generalize hitherto obtained results for multiplayer hypermatrix games. Finally, we conclude in Sec. VII.
II Equivalence between various definitions of ESS in bimatrix Games
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS). As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
II.1 Mutants in both subpopulations
Let us begin with following set of three definitions (aptly termed and-definition) which can be shown to be equivalent:
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{and}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
    Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (3)
    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then for all $\mathbf{{x}}\neq\mathbf{\hat{x}}$,
    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (4)
    This reminds one of what Hofbauer and Sigmund termed the “weak ESS” Hofbauer and Sigmund (1998); it also includes the definitions Selten Selten (1980) and Schuster et al. Schuster et al. (1981); Hofbauer and Sigmund (1988) proposed as ESS for bimatrix games.
  3. Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
    This way of defining ESS proves useful when linking the ESS definition with its dynamical stability. We emphasize that this definition is different from the previous two because it compares the fitnesses of resident states with that of only their nearby states.
It is of use to recall that a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a Nash equilibrium Nash (1950, 1951); Hofbauer and Sigmund (1998) (NE) in bimatrix games if the following is true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$:
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}~{}~{}\text{and}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
Mere inspection reveals that NE is directly implied by and-definition; and, of course, when NE is strict, ESS under and-definition is implied.
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three equivalent definitions (aptly termed sum-definition):
  1. Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (7)
  2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which equality in (i) holds, then
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (9)
  3. Definition 2c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (10)
It is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true. We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
The equilibrium condition of Definition 2b (Eq. (8)) appears different from Eq. (6). So while strict NE obviously implies ESS under sum-definition, it is not obvious if the ESS implies NE always. To check this out, let’s write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$. Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which means both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds good. Concluding, an ESS under sum-definition implies NE.
We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly. In other words, states $(\mathbf{\hat{x}},\mathbf{y})$ and $(\mathbf{x},\mathbf{\hat{y}})$ are not mathematically allowed mutant states. This restriction is relaxed in the next definition as introduced in the next section.
II.2 Mutants in both or either subpopulations
Few remarks are in order. ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants. In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in sum-definition. In the light of above remarks, we now present the following set of equivalent definitions (aptly termed or-definition) where one of $\epsilon_{x}$ or $\epsilon_{y}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
  1. Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$ with $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{x},\epsilon_{y})\neq(0,0)$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{or}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
  2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both the equalities in (i) hold,
    $\text{if}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{then}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}~{}\text{then}~{}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (13)
    (It basically captures the fact that both the mutants cannot simultaneously take advantage of the resident; at least one of them gets penalized.)
  3. Definition 3c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}~{}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998, 2003) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
We notice that (6) directly matches with equilibrium conditions in Definition 3b. So obviously, ESS under or-definition implies NE; and strict NE trivially implies ESS. However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’. However, this apparent anomaly is prevented since the definition allows for one of $\epsilon_{x}$ or $\epsilon_{y}$ to be zero independently. Specifically, in the case pointed out, mathematically setting $\epsilon_{x}=0$ (and $\epsilon_{y}\neq 0$), physically means that no mutants in X-subpopulation are allowed, i.e., no ${\bf x}\neq{\bf{\hat{x}}}$ can be used to satisfy $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, and hence the first inequality in (11) for infinitesimally small $\epsilon_{y}$. Owing to the conjunction ‘or’, the second inequality in (11) must be checked which anyway can’t be satisfied as we have assumed that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. In conclusion, allowing $\epsilon_{x}$ and $\epsilon_{y}$ to vanish (but not simultaneously) prevents states that are not NE’s from qualifying as ESS’s. Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
Consider Definition 3a version of or-definition. If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$. Consequently, setting any of the $\epsilon_{x}$ or $\epsilon_{y}$ to zero is not permissible, as it renders both conditions meaningless. For instance, if we set $\epsilon_{y}=0$, then both the comparisons $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid because $(\epsilon_{x},0)$ physically means there is no mutant (no choice of ${\bf y})$ in the Y-population. Thus, if you would want to allow mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS mutants have to be present in both X- and Y-subpopulations—this perspective may appositely be called the simultaneous deviation perspective. This view is adopted while defining the Nash–Pareto pair in Hofbuaer and Sigmund Hofbauer and Sigmund (1998). The alternate perspective—which we term face-value perspective is to permit all deviations outlined explicitly by the or-definition [as specified in Definition 3a, $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE. This view was adopted by Cressman Cressman and Tao (2014); Cressman (1992) while defining 2ESS. It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
II.3 Equivalent definitions
As mentioned earlier each of aforementioned three sets (viz., and-, sum- and or-definitions) has three equivalent definitions—versions (a), (b) and (c). Let us see the line of reasoning that establishes these equivalences. For illustrative purpose, let us focus on establishing the equivalence between Definitions 3a, 3b, and 3c; equivalence between Definitions 1a, 1b, and 1c, as well as the equivalence between Definitions 2a, 2b, and 2c, follows analogously. Now, in what follows, we prove the assertion by demonstrating that Definitions 3a and 3b imply each other and Definitions 3a and 3c imply each other.
Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all infinitesimal $(\epsilon_{x},\epsilon_{y})\neq(0,0)$, inequalities in (11) are valid. First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there always is some permissible value of $(\epsilon_{x},\epsilon_{y})$ for which the inequalities in (11) do not hold. For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $\epsilon_{x}\neq 0$ and $\epsilon_{y}=0$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid because there is no mutant in Y-population hence the comparison in the second condition in (11) won’t make any physical sense. Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—across the allowable ranges of $\epsilon_{x}$ and $\epsilon_{y}$ to satisfy the inequalities in (11) is:
$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
Thus, four explicit possibilities arise from (15) as detailed below. We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition in (12)) must be satisfied with the natural requirement of conditions in (13) in the relevant case.
  1. 1.
    One can trivially see that either condition— $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$—is true when (11) holds good, or
  2. 2.
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
  3. 3.
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
  4. 4.
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$, straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the conditions in (13) must follow.
Above argument establishes that Definition 3a implies Definition 3b. Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent. We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESS’s possible are the ones that are strict NE.
Next let us show that Definitions 3a and 3c imply each other. Starting with Eq. (11), we multiply both sides of the first inequality of Eq. (11) by $\epsilon_{x}$ and add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides. Similarly, we multiply both sides of the second inequality of Eq. (11) by $\epsilon_{y}$ and add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides. The two resultant inequality expressions can be rearranged to read:
$\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ (16a)
[…]
We define $\mathbf{x_{1}}\equiv\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small. Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq. (14) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
It, however, is interesting to note that and-, sum-, and or-definitions are not equivalent and need not imply each other except that the and-definition implies the sum-definition for any ESS. This aspect of the definitions will be taken up in the next two sections.
III Bimatrix games
Among the three sets of definitions which one is the most useful one? Or each of definitions works equally well? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable, that is, we expect the following claim to be satisfied an ESS Cressman and Tao (2014); Cressman (1996):
Theorem 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
$\dot{x}_{i}=x_{i}\left(({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (17a)
$\dot{y}_{i}=y_{i}\left(({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}\right).$ (17b)
It is well-known that NE corresponds to a fixed point of the replicator equation (17). Since we have already seen that under and-, sum-, and or-definitions ESS is NE, we effectively are sure that any ESS is a fixed point of the replicator equation.
We can categorize ESS into two types: pure ESS and mixed ESS. Pure ESS can represent either a strict or weak NE, while mixed ESS can only correspond to a weak NE. The nuances of stability arise specifically in cases where ESS is a weak pure NE. We will examine each category separately for bimatrix and asymmetric games. However, before proceeding, let us first identify which definition aligns with the defining concept of strong stability Hofbauer and Sigmund (1998, 2003).
III.1 Strong Stability
We note that while $x_{i}$ is the frequency of $i$th type in population X and $y_{i}$ is the frequency of $i$th type in population Y, we never explicitly specified what the corresponding strategies are for the two types. So let us bring it into consideration. So, let $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\cdots,x_{m}\}(\equiv\mathbf{x}$) be, respectively, mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}\}$; similarly, we have $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\cdots,y_{n}\}$($\equiv\mathbf{y}$) be mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\cdots,\mathbf{q}_{n}\}$, respectively. In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted as ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices. The size of ${\sf U}_{1}$ is $N_{1}\times N_{2}$, while ${\sf U}_{2}$ has dimensions $N_{2}\times N_{1}$. In these notations, it is obvious that the matrices, ${\sf B}$ and ${\sf C}$, defined in the previous section are the ones with elements $b_{ij}=\mathbf{p}_{i}\cdot{\sf U}_{1}\mathbf{q}_{j}$ and $c_{ij}=\mathbf{q}_{i}\cdot{\sf U}_{2}\mathbf{p}_{j}$, respectively.
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. In any case, we would say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
To this end, consider general $2\times 2$ payoff matrices:
${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (18)
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b. Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
Now, to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is best presented as
$\dot{x}=x(1-x)[y(\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1})-(1-y)(\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2})],$ (19a)
$\dot{y}=y(1-y)[x(\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1})-(1-x)(\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2})].$ (19b)
We now effectively have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This, in fact, is the case can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$. Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE. However, (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ does not ensure that all the trajectories starting in the neighbourhood of $(0,0)$ would flow towards $(0,0)$—some flows to other phase points on the line $x=0$. Hence, local asymptotic stability of $(0,0)$ is not guaranteed; the fixed point is at best Lyapunov stable, a weaker form of stability. Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, aptly also known as weak ESS Hofbauer and Sigmund (1998).
In summary, ESS under any of the three definitions is strongly stable if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$. However, aforementioned weak stability can only be guaranteed by the and-definition.
III.2 Pure ESS
The pure ESS can be either strict NE or non-strict NE. Let’s first discuss what happens when it is a strict NE.
III.2.1 Strict NE as ESS
One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that satisfying any of these definitions guarantees the local asymptotic stability of the corresponding fixed point. However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition. The argument presented in Section IV.1 can be easily adapted to the case of bimatrix games.
Here, because it is analytically doable for bimatrix games, let us go beyond linear stability analysis for a pure ESS $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ that happens to be an isolated fixed point as well. Consider the following function:
$V_{1}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right).$ (20)
Realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If we additionally can show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point. To this end, on taking the time derivative, we get,
$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}),$ (21)
where we have used Eq. (17). Of course, $\dot{V_{1}}<0$ when either inequalities in (5) or inequality in (10) is satisfied. Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable. We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternate means (see Cressman (1996)).
III.2.2 Non-strict pure NE as ESS
Figure 1: Phase portraits supplementing Examples 1–6 of the main text which should be referred for further details. The red circular marker indicates the ESS (fixed point) under focus. The blue lines are trajectories under replicator equation whereas the green dashed lines are trajectories under adjusted replicator equation.
The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated. For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument would not applicable since the fixed point is not isolated. Below we highlight such points through simple examples.
Example 1 (see Fig. 1a): ESS under sum-definition need not be an ESS under and-definition.
Consider the following general class of bimatrix games:
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (22)
where $b_{1}>0$ and $c_{3}>0$ such that it has a non-strict pure NE, $(\hat{x},\hat{y})=(1,1)$. To confirm whether it is an ESS, we denote the probability vector corresponding to the NE as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$, and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative violating the condition of and-definition. To find whether it satisfies sum-definition of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$ which being positive for infinitesimal $\epsilon$ implies $(\hat{x},\hat{y})$ satisfies sum-definition. In passing, one can check that, as has been generally established in Section II.2, it is 2ESS only under simultaneous deviation perspective.
Next, to ascertain the nature of dynamical stability of the ESS, we write the corresponding replicator equation for payoff matrices in (22):
$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (23)
$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (24)
We can see that the fixed point $(1,1)$ can not be asymptotically stable because $\dot{y}$—being non-positive in the fixed point’s neighbourhood—stops $y$-value of any trajectory starting in the neighbourhood from approaching $1$. This demonstrates that, in the case of non-strict NE, Definition 1c and 3c (face-value perspective) are better indicator of stability than Definition 2c. This is because failing to satisfy Definition 1c or 3c, renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable in the light of Theorem 1.
Example 2 (see Fig. 1b): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
Consider the following example:
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (25)
where $b_{2}>b_{3}$ and $c_{2}>c_{3}$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict NE and we can again take neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=(b_{2}-b_{3})\epsilon\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(c_{2}-c_{3})\epsilon\delta$ which are positive for infinitesimally small $\epsilon$ and $\delta$. Therefore, the NE satisfies and-definition but it does not satisfy or-definition in face-value perspective (since for $\epsilon=0$ both the terms become zero). However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition.
To analyze the dynamical stability of the ESS, we cast the corresponding replicator equations as:
$\dot{x}$ $=(b_{2}-b_{3})x(1-x)(1-y),$ (26)
$\dot{y}$ $=(c_{2}-c_{3})y(1-y)(1-x).$ (27)
It is evident that both $\dot{x}$ and $\dot{y}$ are positive, indicating that the dynamics is driven away from $x=0$ and $y=0$ by any infinitesimal perturbation. Note that since $x=1$ and $y=1$ are line of fixed points, once a trajectory reaches at either boundary, $x=1$ or $y=1$, it terminates. Furthermore, since $\frac{dy}{dx}=\frac{ky}{x}$, where $k\equiv\frac{c_{2}-c_{3}}{b_{2}-b_{3}}$, one arrives at $y=Kx^{k}$ ($K$ being an integration constant) as equation of trajectories. This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$. This implies absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$. Concluding, the non-strict NE as ESS need not satisfy Theorem 1.
III.3 Mixed ESS
The state-strategy correspondence—discussed in Section III.1—in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
$\hat{x}(\hat{y}b_{11}+(1-\hat{y})b_{12})+(1-\hat{x})(\hat{y}b_{21}+(1-\hat{y})b_{22})=x(\hat{y}b_{11}+(1-\hat{y})b_{12})+(1-x)(\hat{y}b_{21}+(1-\hat{y})b_{22}).$ (28)
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, such correspondence is there for Definitions 2b and 3b as well.
Now we recall the well-known result Schuster et al. (1981); Selten (1980) that ESS of Definitions 1c or 2c can never be mixed. There may, however, be mixed 2ESS—the ESS of Definition 3c (as shown below) under simultaneous deviation perspective. Since 2ESS must be NE which must be a fixed point of the replicator equation, let us assume that it corresponds to a fixed point $(\hat{x},\hat{y})$. A glance at replicator equation (19) reveals that with $\alpha_{1}\equiv\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$,  $\beta_{1}\equiv\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$,  $\alpha_{2}\equiv\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and $\beta_{2}\equiv\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$, the internal fixed point must be $(\hat{x},\hat{y})=\left({\alpha_{2}}/({\alpha_{2}+\beta_{2}}),{\alpha_{1}}/({\alpha_{1}+\beta_{1}})\right)$. The eigenvalues of Jacobian in linearization about this fixed points can be calculated to be ±$\pm\sqrt{({\alpha_{1}\beta_{1}\alpha_{2}\beta_{2}})/[{(\alpha_{1}+\beta_{1})(\alpha_{2}+\beta_{2})}]}$. Evidently, the fixed point can either be a center or a saddle. We know that (linear) center found using the linear stability analysis may not actually be a robust nonlinear center. However, it is known that the system is Hamiltonian Hofbauer and Sigmund (1998) in the interior of the phase space, and hence any center in this case is, in fact, nonlinear center. Thus, ESS of any definition is not strongly stable.
In conclusion, we find that Theorem 1 is invalid if the 2ESS is mixed. The theorem, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
$\dot{x}_{i}=x_{i}\frac{({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}}{\mathbf{x}\cdot{\sf B}\mathbf{y}},$ (29a)
$\dot{y}_{i}=y_{i}\frac{({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}}{\mathbf{y}\cdot{\sf C}\mathbf{x}}.$ (29b)
Under its dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1c).
Example 3 (see Fig. 1c): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
To simplify the calculation while keeping generality, we use two key facts: first, one can always modify the payoff matrix by adding a matrix with identical rows and by scaling the payoff matrices, which preserves both the dynamics and the ESS. Specifically, ${\sf B}$ and $c{\sf B}+{\sf\Omega}$ are equivalent, where $c$ is a positive scalar and ${\sf\Omega}_{ij}=\omega_{j}$. Second, using this transformation, one can show that any $2\times 2$ bimatrix game can be reduced to either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). Therefore, we can work with the following payoff matrix:
${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}~{}\text{and}~{}{\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (30)
where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998). Now, the internal fixed point—if it exist in the simplotope—is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2}),{\beta_{2}}/({\alpha_{2}+\beta_{2}})\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}}),{\beta_{1}}/({\alpha_{1}+\beta_{1}})\end{pmatrix}\right)$ and the neighbourhood vector is $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2})+\delta_{x},{\beta_{2}}/({\alpha_{2}+\beta_{2}})-\delta_{x}\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}})+\delta_{y},{\beta_{1}}/({\alpha_{1}+\beta_{1}})-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small which can have positive or negative sign but neither of them are allowed to be zero as per simultaneous deviation perspective. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=-\delta_{x}\delta_{y}(\alpha_{1}+\beta_{1})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta_{x}\delta_{y}(\alpha_{2}+\beta_{2})$. We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game. The fixed point is always an ESS under or-definition if the they have different sign, this is the case of rescaled zero-sum game. Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
Therefore, we conclude that mixed ESS under or-definition is possible, however it does not tell anything about asymptotic stability—an ESS can at best be Lyapunov stable as a center. This also tells us that unlike the case of strict NE, and-, sum- and or-definitions are not equivalent when it comes to mixed NE; in fact there is no scope of mixed ESS under and- and sum-definitions. Finally, of course, as per the face-value perspective the interior fixed point is not an ESS, since $\delta_{x}=0$ makes $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=0$.
IV Asymmetric games
It is now natural to ask if introducing some intra-specific interactions changes this situation, i.e., renders Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games). In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games (see also Appendix A).
Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intra-specific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
  1. Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{\emph{and}}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (31)
    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s. Mathematically, we may write,
  2. Definition 2d (sum-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (32)
    As discussed in Section II, since the above definitions lead to only pure ESS, Cressman Cressman (1992, 1996) gave a weaker definition of ESS for asymmetric games. It says that, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous. Mathematically,
  3. Definition 3d (or-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{\emph{or}}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (33)
    ESS under this definition is also called 2ESS as before. When both inter- and intra-specific interactions are simultaneously present, mixed 2ESS may appear even without the imposing of additional simultaneous deviation perspective, which thus is no longer adopted henceforth.
Let us also, for clarity, reiterate Theorem 1 in the context of asymmetric games Cressman (1992, 1996).
  1. Theorem ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
    $\dot{x}_{i}=x_{i}\left(({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (34a)
    $\dot{y}_{j}=y_{j}\left(({\sf C}\mathbf{x})_{j}+({\sf D}\mathbf{y})_{j}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}\right).$ (34b)
We now turn to discussing pure and mixed ESS’s separately in the context of asymmetric games.
IV.1 Pure ESS
In case of a strict NE being ESS, all the eigenvalues of the Jacobian about the corresponding fixed point of the replicator dynamics are negative making the fixed point locally asymptotically stable under all the three definitions (Definitions 1d, 2d, and 3d). To demonstrate this, we go beyond the restriction of two strategies per subpopulation and address the matter in its full generality. We remind ourselves that these definitions are valid for any finite number of strategies, i.e., for arbitrary $m$ and $n$ values. In fact—as mentioned in Section III.2—even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
IV.1.1 Strict NE as ESS
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (34a) and (34b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity. It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
Now we would like to figure out whether the strict NE is ESS. For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative. A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\delta$’s:
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$ (35a)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}},$ (35b)
$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\delta_{x_{1}}+(a_{2m}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\delta_{x_{m-1}}$ (35c)
$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$
$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}}$ (35d)
$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\delta_{x_{1}}+(b_{2n}-b_{mn})\delta_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\delta_{x_{m-1}}.$
Therefore, recalling the forms of the eigenvalues, we find that
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\delta_{x_{1}}\lambda_{1}-\delta_{x_{2}}\lambda_{2}-\cdots-\delta_{x_{m-1}}\lambda_{m-1}.$ (36)
Similarly, we can calculate
$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\delta_{y_{1}}\lambda^{\prime}_{1}-\delta_{y_{2}}\lambda^{\prime}_{2}-\cdots-\delta_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (37)
We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq. (36). Similarly, last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq. (37). If the fixed point has to be ESS then for both equations, the RHS’s have to be positive in the entire neighbourhood. This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero. This demand is automatically satisfied if the fixed point is strict NE.
We furthermore find that if we add Eq. (36) and Eq. (37) we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}.$ (38)
The LHS has to be positive for the fixed point to be ESS under sum-definition. It is easy to see that exactly same argument as given above for the ESS under and-definition applies to this case. Hence, in the context of linear stability, and- and sum-definitions appear equivalent.
Lastly, in the light of Eq. (36) and Eq. (37), or-definition implies
either $\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (39)
for the fixed point to be ESS. The validity of condition (39) in the entire neighbourhood requires the negativity of all the eigenvalues, or equivalently, requires the fixed point to be strict NE. It might be tempting to assume that not all eigenvalues need to be negative for condition (39) to hold, given the presence of conjunction or. However, it is important to recall that in Definition 3d, mutants in either or both subpopulations are allowed, as discussed in Section II.2. Therefore it is allowed to have either all $\delta_{x}$’s or all $\delta_{y}$’s to be zero rendering either the former or the latter inequality in condition (39) invalid. As a result, the inequalities in (39) imply that all eigenvalues must be negative. In conclusion, strict NE is 2ESS as well.
IV.1.2 Non-strict NE as ESS
The equivalence between ESS’s under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
Example 4 (see Fig. 1d): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
Consider
${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (40)
where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$. We can see that $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is a non-strict NE whose neighbourhood state is given as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. To check whether it is an ESS we calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon\delta(b_{2}-b_{3})+\epsilon^{2}(a_{2}-a_{3})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon\delta(c_{2}-c_{3})+\delta^{2}(d_{2}-d_{3})$. It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$. However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy an ESS condition under sum-definition. The fixed point remains asymptotically stable as may be checked numerically, thus, validating Theorem $1^{\prime}$.
Before going to the case of mixed ESS, we would like to bring it to the fore that for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Theorem 1 in case of pure ESS corresponds to an isolated fixed point. One can show using the same Lyapunov function given in Eq. (20), that satisfying Definition 1d or 2d imply asymptotic stability for the pure ESS (corner fixed point of the simplotope). However, the general proof that Definition 3d imply asymptotic stability for pure ESS is not so straightforward: A proof using center manifold theory is available in the literature Cressman (1996).
IV.2 Mixed ESS
The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games. Specifically now and- and sum-definitions can lead to mixed ESS and mixed ESS under or-definition can be locally asymptotically stable. While the former is illustrated in Example 5 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, it it exists, is
$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right).$ (41)
Here, $a\equiv a_{11}-a_{12}-a_{21}+a_{22}$$a^{\prime}\equiv a_{12}-a_{22}$$b\equiv b_{11}-b_{12}-b_{21}+b_{22}$$b^{\prime}\equiv b_{12}-b_{22}$$c\equiv c_{11}-c_{12}-c_{21}+c_{22}$$c^{\prime}\equiv c_{12}-c_{22}$$d\equiv d_{11}-d_{12}-d_{21}+d_{22}$ and $d^{\prime}=d_{12}-d_{22}$. On linearization, the eigenvalues, say, $\lambda_{-}$ and $\lambda_{+}$, of the Jacobian matrix comes out as
$\lambda_{\pm}=\frac{(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))\pm i\sqrt{4\hat{x}(1-\hat{x})\hat{y}(1-\hat{y})(ad-cb)-(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))^{2}}}{2}.$ (42)
It is evident that now there is a possibility of focus or node types of fixed points under suitable conditions. For clear visualization, consider the example of asymmetric game given in (44), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig. 1(c) and Fig. 1(f). Furthermore, Example 6 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
Example 5 (see Fig. 1e): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
Suppose that
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (43)
One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$. Therefore the neighbourhood vector has the form—$(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}{2}/{9}+\delta_{x},{7}/{9}-\delta_{x}\end{pmatrix},\begin{pmatrix}{2}/{9}+\delta_{y},{7}/{9}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small and can be positive or negative but they can’t be zero simultaneously. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}={3\delta_{x}^{2}}/{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}={3\delta_{y}^{2}}/{4}$. Both terms are positive for all infinitesimal values of $\delta_{x}$ and $\delta_{y}$, therefore it satisfies and-definition. For sum-definition one just need to add both the terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}={3(\delta_{x}^{2}+\delta_{y}^{2})}/{4}$, which means sum-definition is also satisfied. Fig. 1(e) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Theorem 1 for and- and sum-definitions.
Example 6 (see Fig. 1f): Mixed 2ESS need not be ESS under and- and sum-definitions.
Consider the payoff matrices,
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (44)
Now, let see whether the internal fixed point is an ESS. One can calculate the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}66/115,49/115\end{pmatrix},\begin{pmatrix}238/345,107/345\end{pmatrix}\right)$, the neighbourhood vector therefore is, $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\frac{66}{115}+\delta_{x},\frac{49}{115}-\delta_{x}\end{pmatrix},\begin{pmatrix}\frac{238}{345}+\delta_{y},\frac{238}{345}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small. We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=3\delta_{x}\delta_{y}+\frac{3\delta_{x}^{2}}{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-7\delta_{x}\delta_{y}+\frac{3\delta_{y}^{2}}{4}$. It is not very hard observe that at least one of the terms in the RHS of the two expressions is positive for all possible infinitesimal values of $\delta_{x}$ and $\delta_{y}$, which means the internal fixed point is 2ESS—an ESS under or-definition. Since both the terms cannot remain simultaneously positive in the entire neighbourhood, it is definitely not an ESS under and-definition. To check for sum-definition we add the two terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-4\delta_{x}\delta_{y}+\frac{3(\delta_{x}^{2}+\delta_{y}^{2})}{4}$, which is not guaranteed to be always positive in the neighbourhood; hence it is not an ESS under sum-definition as well. Fig. 1(f) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Theorem 1 for or-definition.
The stability of the mixed ESS can be understood through Lyapunov function. Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary by considering the following function given in Eq. (20). We calculate the time derivative using Eq. (34) and get,
$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}),$ (45)
We can see that $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})$ in some neighbourhood of $({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (31) or inequalities (32) are satisfied. Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. Note that, these conclusions does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) for mixed ESS since we know only pure ESS is possible there. As far as the or-definition is concerned, a theorem Cressman (1996) comes to our help:
  1. Theorem 2: $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a mixed 2ESS if and only if there exists a positive $r$ (termed ‘weight factor’) such that $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}+r(\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x})>0.$ (46)
In other words, this condition is an equivalent qualifying condition for 2ESS defined in Definition 3d. Consequently, following similar steps given earlier, it is now straightforward to see that
$V_{r}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+r\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right),$ (47)
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
V Information theoretical aspects
The mention of KL-divergence (also, known as relative entropy) in the preceding sections is quite intriguing. The fact that a core concept of information theory is tied to the stability notion of ESS, reminds us of a known result in symmetric games in line with the general principle of minimum relative entropy Bomze (1991); Karev (2010); Qian et al. (1991); Floerchinger and Haas (2020): The KL-divergence of ESS from another time-evolving state monotonically decreases in symmetric games Baez and Pollard (2016). In asymmetric games, however, the presence of multiple definitions of ESS (the and-, sum-, and or-definitions) necessitates different adaptations of the KL-divergence to establish the relationship between KL-divergence and ESS. Furthermore, since Theorem 2 applies only to mixed states, the information-theoretic connection of the or-definition is also valid solely for mixed states.
For completeness, let us recall that mathematically KL-divergence of a probability distribution with probability mass function (pmf) $p$ from another probability distribution ($q$ being the pmf) defined over the same random variable is expressed as
$D_{\rm KL}({p}||{q})=\sum_{i}p_{i}\log\left[\frac{p_{i}}{q_{i}}\right],$ (48)
where $i$ belongs to the support of the pmf’s $p$ and $q$. The KL-divergence is a useful tool for quantifying the dissimilarity (or distance) between two probability distributions. In the context of population dynamics, it may be employed to measure the difference between two population states as they are nothing but frequencies which qualify as probabilities Chakraborty (2024).
What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions. However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with and-definition: It is obvious that the fixed point corresponding to ESS of and-definition, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
We find the weight factor, $r$, interesting and would like to find some game-theoretic connection for it. Below, we establish such a connection at least for a specific category of games, which we call asymmetric partnership game. In an asymmetric partnership game, the symmetric component mirrors a partnership game (i.e., ${\sf A}^{T}={\sf A}$ and ${\sf D}^{T}={\sf D}$) and the bimatrix component of the asymmetric game mirrors a rescaled partnership game Hofbauer and Sigmund (1998), i.e., a game with ${\sf B}={\sf G}+{\sf\Omega}$ and ${\sf C}=c{\sf G}^{T}+{\sf\Omega^{\prime}}$ for some ${\sf G}$ and $c>0$, where ${\sf\Omega}$ and ${\sf\Omega^{\prime}}$ matrices are such ${\sf\Omega}_{ij}=\omega_{j}$ and ${\sf\Omega}^{\prime}_{ij}=\omega^{\prime}_{j}$, with $\omega_{j},\omega^{\prime}_{j}\in\mathbb{R}$. We intend to show that $r=c^{-1}$ in asymmetric partnership game. We accomplish this in two steps: In the first step, we determine the potential function whose Shahshahani gradient corresponds to asymmetric replicator dynamics; in the second step, we construct a Lyapunov function from this potential and demonstrate that the time derivative of this Lyapunov function is a negative quantity, implying that the associated fixed point happens to be asymptotically stable.
We begin by putting forward the potential function whose Shahshahani gradient Shahshahani (1979) yields the asymmetric replicator dynamics (Eq. (34)). First, we define an inner product in the tangent space:
$\langle(\bm{\xi}^{\prime},\bm{\eta}^{\prime}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\frac{\xi_{i}^{\prime}\xi_{i}}{x_{i}}+\frac{1}{c}\sum_{i=1}^{n}\frac{\eta_{i}^{\prime}\eta_{i}}{y_{i}},$ (49)
where $\bm{\xi}^{\prime}$ and $\bm{\xi}$ are tangent vectors tangent to $\Sigma_{m}$ whereas $\bm{\eta}^{\prime}$ and $\bm{\eta}$ are tangent vectors tangent to $\Sigma_{n}$. To write the metric involved in the inner product explicitly, we define $\bm{\zeta}^{\prime}=(\bm{\xi}^{\prime},\bm{\eta}^{\prime})$  and  $\bm{\zeta}=(\bm{\xi},\bm{\eta})$, where the first $m$ components are $x$-components and form $m+1$ to $m+n$ are $y$-components. Under the coordinate transformation, $z_{i}=x_{i}$ for $i$ from $1$ to $m$ and $z_{i}=cy_{i-m}$ for $i$ from $m+1$ to $n+m$, we may express the metric as,
$g_{ij}(z)=\frac{\delta_{ij}}{z_{i}},$ (50)
which is the Shahshahani metric. Moreover, we can recast Eq. (49) as,
$\langle\bm{\zeta}^{\prime},\bm{\zeta}\rangle=\sum_{i=1}^{m+n}\frac{\zeta^{\prime}_{i}\zeta_{i}}{z_{i}}=\sum_{i,j=1}^{m+n}g_{ij}\zeta^{\prime}_{i}\zeta_{j}.$ (51)
We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))\to(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
${\sf J}_{ij}\equiv\mathbb{E}\left[\frac{\partial\log p}{\partial x_{i}}\frac{\partial\log p}{\partial x_{j}}\right]\to\mathbb{E}\left[\frac{\partial\log x}{\partial x_{i}}\frac{\partial\log x}{\partial x_{j}}\right]=\sum_{k}x_{k}\frac{\delta_{ik}}{x_{k}}\frac{\delta_{jk}}{x_{j}}=\frac{\delta_{ij}}{x_{i}}=g_{ij}.$ (52)
This relation is true even in asymmetric games, where under a similar mapping, $1/(1+c)$-scaled Fisher information metric becomes Shahshahani metric.
Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is replicator equation):
$\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\frac{\partial U}{\partial x_{i}}\xi_{i}+\sum_{i=1}^{n}\frac{\partial U}{\partial y_{i}}\eta_{i}.$ (53)
We calculate the L.H.S. of this equation, using Eq. (49) and replicator equation, Eq. (34):
$\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\xi_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}]+\frac{1}{c}\sum_{i=1}^{n}\eta_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}],$ (54)
since $\bm{\xi}$ and $\bm{\eta}$ being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively. As far as the R.H.S. of Eq. (54) is concerned, one can easily verify that following is the potential whose Shahshahani gradient is replicator dynamics by substituting it in the R.H.S. of Eq. (53):
$U\equiv\frac{1}{2}\left[\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x}\right].$ (55)
To arrive at this conclusion, we must invoke the properties of ${\sf\Omega}$ and ${\sf\Omega}^{\prime}$ that $\sum_{i}\xi_{i}({\sf\Omega}\mathbf{y})_{i}=0$ and $\sum_{i}\eta_{i}({\sf\Omega}^{\prime}\mathbf{x})_{i}=0$ owing to the respective facts that $({\sf\Omega}\mathbf{y})_{i}$ and $({\sf\Omega}^{\prime}\mathbf{x})_{i}$ are independent of the index $i$.
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value—in fact, 00. Now, extrema of $V$ is attained at a point where
$\frac{\partial V}{\partial x_{i}}=0\implies({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}=\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y},$ (56a)
$\frac{\partial V}{\partial y_{i}}=0\implies({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}=\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (56b)
This means that the extrema $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics, Eq. (34). For this fixed point to be minima, the Hessian matrix,
${\sf H}=\begin{bmatrix}\left[\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}\right]_{m-1,m-1}&\left[\frac{\partial^{2}V}{\partial x_{i}\partial y_{j}}\right]_{m-1,n-1}\\ \left[\frac{\partial^{2}V}{\partial y_{i}\partial x_{j}}\right]_{n-1,m-1}&\left[\frac{\partial^{2}V}{\partial y_{i}\partial y_{j}}\right]_{n-1,n-1}\end{bmatrix}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-\frac{1}{c}(c_{ij}-c_{im}-c_{nj}+c_{nm})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}$ (57)
has to be positive definite. Here, as before, we have used $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s as elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Using the relation between ${\sf B}$ and ${\sf C}$ (i.e., ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$), one can show that,
${\sf H}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-(b_{ji}-b_{mi}-b_{jn}+b_{mn})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}.$ (58)
We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minima Kuttler (2008). Supposing a minimum exists at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, we evaluate the time derivative of $V$ to be
$\dot{V}=-\sum_{i=1}^{m}\dot{x}_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}]-\frac{1}{c}\sum_{i=1}^{n}\dot{y}_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}].$ (59)
Here we have used the following to get Eq. (59): $\frac{d}{dt}(\mathbf{x}\cdot{\sf A}\mathbf{x})=2\sum_{i}\dot{x}_{i}({\sf A}\mathbf{x})_{i}$, $\frac{d}{dt}(\mathbf{x}\cdot{\sf B}\mathbf{y})=\sum_{i}\dot{x}_{i}({\sf B}\mathbf{y})_{i}+\frac{1}{c}\sum_{i}\dot{y}_{i}({\sf C}\mathbf{x})_{i}+\sum_{i}\dot{y}_{i}({\Omega}^{T}\mathbf{x})_{i}$$\frac{d}{dt}(\frac{\mathbf{y}\cdot{\sf C}\mathbf{x}}{c})=\frac{1}{c}\sum_{i}\dot{y}_{i}({\sf C}\mathbf{x})_{i}+\sum_{i}\dot{x}_{i}({\sf B}\mathbf{y})_{i}+\frac{1}{c}\sum_{i}\dot{x}_{i}({\sf\Omega^{\prime}}^{T}\mathbf{y})_{i}$ and $\frac{d}{dt}(\mathbf{y}\cdot{\sf D}\mathbf{y})=2\sum_{i}\dot{y}_{i}({\sf D}\mathbf{y})_{i}$. Furthermore, using the replicator equation Eq. (34), we get,
$\frac{dV}{dt}=-\sum_{i=1}^{m}x_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}]^{2}-\frac{1}{c}\sum_{i=1}^{n}y_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}]^{2}.$ (60)
This implies that $\dot{V}<0$ except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $\dot{V}=0$. Consequently, $V$ serves as the Lyapunov function indicating that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is locally asymptotically stable.
Since $V(\mathbf{\hat{x}},\mathbf{\hat{y}})<V(\mathbf{x},\mathbf{y})$, by definition,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{\hat{x}}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}-\mathbf{\hat{x}}\cdot{\sf\Omega}\mathbf{\hat{y}}+\frac{1}{c}\hat{\mathbf{y}}\cdot{\sf D}\mathbf{\hat{y}}+\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}-\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf\Omega^{\prime}}\mathbf{\hat{x}}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x},$ (61)
for all the neighborhood ($\mathbf{x},\mathbf{y}$) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$). Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood Hofbauer and Sigmund (1998) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$), we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}+\frac{1}{c}(\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x})>0.$ (62)
On comparing this expression with Eq. (46), one concludes that $r=1/c$—what we aimed to prove.
VI ESS in Multiplayer hypermatrix games
In the preceding sections, we delved into the definitions of ESS and elucidated their respective domains of applicability, i.e., we clarified which definitions are useful for characterizing pure ESS and which are useful for mixed ESS, all within the context of two-player interactions. However, what if the strategic interaction involves more than two players? In such scenarios, we must shift our focus to the framework of multiplayer games Broom et al. (1997); Broom and Rychtář (2022); Gokhale and Traulsen (2010). Within the realm of multiplayer games, questions might arise, what is the definition of ESS? Are there multiple definitions of ESS, and what are their respective domains of applicability? In this section, we address these questions. Multiplayer games can be somewhat intricate. For the purpose of this paper, it suffices to focus—unless otherwise specified—on multiplayer hypermatrix games where the payoffs are linear in their respective arguments Broom and Rychtář (2022).
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $\mathbf{p}_{i}$-strategied opponent.
In order to get familiarized with the notations to be used in this section, let’s consider the scenario of symmetric hypermatrix games with two strategies, $\mathbf{\hat{p}}$ and $\mathbf{p}$, both belong to some $\Sigma_{N}$, in the population. By definition, the game is symmetrical among all players, meaning that the order of the players is irrelevant and the set of strategies available is the same for all players. In case all the interacting players adopt the same mixed strategy, $\mathbf{\hat{p}}$, the payoff should notationaly be ${\sf U}[\mathbf{\hat{p}};\mathbf{\hat{p}}^{\nu-1}]$. For $\mathbf{\hat{p}}$ to be an Evolutionary Stable Strategy (ESS), it should be advantageous against small mutations of type $\mathbf{p}$. Let $\epsilon$ represents the frequency of players choosing $\mathbf{p}$ and $1-\epsilon$ denotes the frequency of players choosing $\mathbf{\hat{p}}$. In other words, a strategy $\mathbf{\hat{p}}$ is an ESS if for any mutant strategy $\mathbf{p}$$\mathbf{p}\neq\mathbf{\hat{p}}$, there exists an invasion barrier $\epsilon_{\mathbf{p}}>0$, such that $\forall\epsilon$, with $0<\epsilon<\epsilon_{\mathbf{p}}$,
${\sf U}\left[\mathbf{\hat{p}};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right]>{\sf U}\left[\mathbf{p};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right].$ (63)
Since the payoff is linear in arguments, we recognize that
${\sf U}\left[\mathbf{\hat{p}};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right]=\sum_{k=0}^{\nu-1}\binom{\nu-1}{k}\epsilon^{k}(1-\epsilon)^{\nu-k-1}{\sf U}[\mathbf{\hat{p}};\mathbf{p}^{k},\mathbf{\hat{p}}^{\nu-k-1}].$ (64)
Introducing a short hand notation for average population strategy as $\tilde{\mathbf{p}}=\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}}$, inequality (63) can be rewritten as, ${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$. (Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (63) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
$\epsilon{\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>\epsilon{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (65)
Absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
${\sf U}\left[\epsilon\mathbf{\hat{p}}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (66)
Thus, we arrive at the final conclusion that a strategy $\mathbf{\hat{p}}$ is an ESS if for every strategy $\tilde{\mathbf{p}}$ that are sufficiently close but not equal to $\mathbf{\hat{p}}$,
${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\tilde{\mathbf{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (67)
This definition is analogous to Definitions 1c, 1d, 2c, 2d, 3c, and 3d of ESS in the case of two-player games.
However, note that those definitions are in terms of the states and not strategies. When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of $k$-th strategy. Here, we have adopted a notational convention: When $\mathbf{p}_{i}$ appears inside square brackets, as in ${\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]$, it represents the unit vector in the direction of the $i$-th pure state on the simplex $\Sigma_{m}$, rather than the strategy $\mathbf{p}_{i}$ itself. The value of ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$ is given by $\sum_{{i_{2}}{i_{3}},\cdots,{i_{\nu}}}x_{i_{2}}x_{i_{3}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Similarly, ${\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}x_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$, and ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}\hat{x}_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[{\mathbf{p}}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Finally, in the context of states we can say that a state $\mathbf{\hat{x}}$ is an Evolutionary Stable State (ESS) if
${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]>{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right].$ (68)
for every $\mathbf{x}\in\Sigma_{m}$ that are sufficiently close but not equal to $\mathbf{\hat{x}}$.
Of course, in the notation above, the replicator dynamics for the symmetric hypermatrix games can be written as,
$\dot{x}_{i}=x_{i}\left({\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right).$ (69)
To explore asymptotic stability of its fixed point, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$. This function is positive definite by definition of KL-divergence and $V_{s}(\mathbf{\hat{x}})=0$. We evaluate
$\dot{V}_{s}=-\sum_{i}\hat{x}_{i}\frac{\dot{x}_{i}}{x_{i}}=-\sum_{i}\hat{x}_{i}\left({\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right)=-\left({\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right).$ (70)
This implies that if ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]>{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]$ (i.e., if $\mathbf{\hat{x}}$ is ESS), then the corresponding fixed point is (locally) asymptotically stable as then $\dot{V}_{s}<0$.
VI.1 Asymmetric hypermatrix games
We are now ready to delve into the definitions of ESS in asymmetric hypermatrix games in a population consisting of two subpopulations. When X-subpopulation is in state $\mathbf{x}$ ($m$-component vector) and Y-subpopulation is in state $\mathbf{y}$ ($n$-component vector), we define the average payoff of a player in X-subpopulation as ${\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]$. Note that a player in X is involved in strategic game with $\nu_{x}-1$ players of X and $\gamma_{x}$ players of Y. Similarly, the average payoff of a Y-player is given as ${\sf C}[\mathbf{y};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]$—a player in Y is involved in strategic game with $\nu_{y}$ players of X and $\gamma_{y}-1$ players of Y. To remain notationaly consistent with the previous sections, we have used ${\sf A}$ and ${\sf D}$ for intra-specific interactions and ${\sf B}$ and ${\sf C}$ for inter-specific interactions. Letting boldfaced letters ${\mathbf{p}}$ and ${\mathbf{q}}$ denote the strategies in X and Y subpopulations, respectively, the replicator equation can be expressed as
$\dot{x}_{i}=x_{i}\left({\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{i};\mathbf{y}^{\gamma_{x}}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]\right),$ (71a)
$\dot{y}_{j}=y_{j}\left({\sf C}[\mathbf{q}_{j};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{j};\mathbf{y}^{\gamma_{y}-1}]-{\sf C}[\mathbf{y};\mathbf{x}^{\nu_{y}}]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right).$ (71b)
Now, one could define three non-equivalent definitions of ESS:
  1. Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{and}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (72)
  2. Definition 2e (sum-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]+{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (73)
  3. Definition 3e (or-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{or}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (74)
    In this case, the ESS may be called 2ESS in keeping with the earlier discussions on two-player games.
The corresponding definitions of ESS for bihypermatrix games can be trivially read off from above three definitions on putting ${\sf A}={\sf D}=0$ in the above definitions.
VI.1.1 Pure ESS
ESS under and-definition for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (72) as
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]=(\hat{x}_{1}-x_{1})\sum_{i_{1},i_{2},\cdots,i_{\gamma_{x}}=1}^{n}y_{i_{1}}y_{i_{2}}\cdots y_{i_{\gamma_{x}}}\left({\sf B}[\mathbf{p}_{1};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]\right)+\cdots$
$+(\hat{x}_{m-1}-x_{m-1})\sum_{i_{1},i_{2},\cdots,i_{\gamma_{x}}=1}^{n}y_{i_{1}}y_{i_{2}}\cdots y_{i_{\gamma_{x}}}\left({\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]\right)>0,$ (75)
and the second inequality in (72) as
${\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]=(\hat{y}_{1}-y_{1})\sum_{i_{1},i_{2},\cdots,i_{\nu_{y}}=1}^{m}x_{i_{1}}x_{i_{2}}\cdots x_{i_{\nu_{y}}}\left({\sf C}[\mathbf{q}_{1};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]\right)+\cdots$
$+(\hat{y}_{n-1}-y_{n-1})\sum_{i_{1},i_{2},\cdots,i_{\nu_{y}}=1}^{m}x_{i_{1}}x_{i_{2}}\cdots x_{i_{\nu_{y}}}\left({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]\right)>0.$ (76)
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under sum-definition (see Appendix B).
Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (71). As far as the pure ESS’s under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (71). Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq. (71) to arrive at
$\frac{dV_{1}}{dt}=-\left({\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]+{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right),$ (77)
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (71) even in the multiplayer case when it is a strict NE. Using the same argument we can also show that all three definitions are equivalent in case of strict NE. We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.
It is not very hard to see that, for ESS that is non-strict pure NE, the situation is more nuanced than for ESS that is strict NE. In fact, similar or even more intricate subtleties may arise, as observed in two-player games.
VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation. The case of 2ESS (or-definition) is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.3). We now show that— extending the argument presented in Ref. Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (71) with ${\sf A}={\sf D}=0$ satisfies the following condition:
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]=\cdots={\sf B}[\mathbf{p}_{m};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}],$ (78a)
${\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]=\cdots={\sf C}[\mathbf{q}_{n};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (78b)
On linearizing the replicator equation about the fixed point, takes the form
${\sf J}=\begin{bmatrix}{\sf 0}&{\sf\alpha}(\hat{x},\hat{y})\\ {\sf\beta}(\hat{x},\hat{y})&{\sf 0}\end{bmatrix},$ (79)
where diagonal elements are zero matrices of sizes $m-1\times m-1$ and $n-1\times n-1$, respectively, while the other two are finite matrices denoted by ${\sf\alpha}$ and ${\sf\beta}$ of sizes $m-1\times n-1$ and $n-1\times m-1$, respectively. The reason for the appearance of zero matrices is the fact that the corresponding elements of Jacobian matrix at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be calculated to be
$\frac{\partial\dot{x}_{i}}{\partial x_{k}}$ $=\frac{\partial}{\partial x_{k}}\left\{x_{i}\left({\sf B}[\mathbf{p}_{i};\mathbf{y}^{\gamma_{x}}]-{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]\right)\right\}=-x_{i}\frac{\partial}{\partial x_{k}}{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]$ (80)
$=-x_{i}\frac{\partial}{\partial x_{k}}\left\{\sum_{j=1}^{m-1}x_{j}{\sf B}[\mathbf{p}_{j};\mathbf{y}^{\gamma_{x}}]+\left(1-\sum_{j=1}^{m-1}x_{j}\right){\sf B}[\mathbf{p}_{m};\mathbf{y}^{\gamma_{x}}]\right\}=-x_{i}\left({\sf B}[\mathbf{p}_{k};\mathbf{\hat{y}}^{\gamma_{x}}]-{\sf B}[\mathbf{p}_{m};\mathbf{\hat{y}}^{\gamma_{x}}]\right)=0.$
Similarly, we can prove that ${\partial\dot{y}_{j}}/{\partial y_{k}}=0$. Now, the characteristic polynomial, $p(\lambda)$ (say), can be obtained as $\det({\sf J}-\lambda I)$. It is evident that $p(\lambda)=(-1)^{(m-1)+(n-1)}p(-\lambda)$—a result derived by changing the signs of the first $m-1$ columns and the last $n-1$ rows of $\det({\sf J}-\lambda I)$. Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also an eigenvalue, thereby preventing the interior fixed point from being sink or source.
Next, we come to mixed 2ESS in asymmetric hypermatrix game. One could expect that as done previously in Sec. IV.2, $V_{r}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ as a candidate for Lyapunov function. For this function to be a valid Lyapunov function one needs an analog of Theorem 2 for multiplayer games, that is, the following proposition needs to be true:
  1. Proposition 1: $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a mixed 2ESS in asymmetric hypermatrix games if and only if there exists an $r>0$ such that $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+r\left({\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right)>0.$ (81)
Of course, it is trivially true that if inequality (81) is satisfied, then $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a 2ESS, meaning that inequalities (74) are also satisfied. However, the converse may not be true making the Proposition 1 false. We present an example below to elucidate this fact. However, it is clear that if a 2ESS satisfies inequality (81), then $V_{r}$ is a right Lyapunov function to chose to prove that 2ESS implies asymptotic stability. The failure to recognize appropriate Lyapunov function doesn’t falsify the statement that 2ESS is asymptotically stable fixed point of the replicator equation. In fact, in our numerical search, we could not find a situation when 2ESS in asymmetric hypermatrix games is not locally asymptotically stable fixed point of the replicator equation. Hence, we propose a conjecture that needs to be proved in due course:
  1. Conjecture 1: A mixed 2ESS in asymmetric hypermatrix games is locally asymptotically stable fixed point of the replicator equation (71).
Let’s consider an insightful example of a 3-player $2\times 2\times 2$ hypermatrix game demonstrating that Proposition 1 is not necessarily valid for all payoff hypermatrices. We consider payoff hypermatrices as shown in Fig. 2.
Figure 2: Figure lists the 3-player game payoff hypermatrices: The four hypermatrices represent the payoffs ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$. The number written at the corners of each of the four hypermatrices are payoff elements. In the first hypermatrix, we have illustrated the convention with which we have assigned the indices of the $2\times 2\times 2$ dimensional hypermatrices.
In order to render calculations analytically tractable, we adopt the assumption that intra-specific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$. Using the payoff hypermatrices of Fig. 2, we can rewrite Eq. (71) as
$\dot{x}=x(1-x)[2y^{2}+y(1-y)-3(1-y)^{2}+\mu\left(-3x^{2}+x(1-x)+(1-x)^{2}\right)],$ (82a)
$\dot{y}=y(1-y)[-2x^{2}-3x(1-x)+5(1-x)^{2}+\mu\left(-3y^{2}+y(1-y)+(1-y)^{2}\right)].$ (82b)
Let us define the following short hand notations: $f_{1}(y)\equiv 2y^{2}+y(1-y)-3(1-y)^{2}$$f_{2}(x)\equiv-3x^{2}+x(1-x)+(1-x)^{2}$$g_{1}(x)\equiv-2x^{2}-3x(1-x)+5(1-x)^{2}$ and $g_{2}(y)\equiv-3y^{2}+y(1-y)+(1-y)^{2}$. For finding interior fixed points, one has to solve $f_{1}(y)+\mu f_{2}(x)=0$ and $g_{1}(x)+\mu g_{2}(y)=0$ simultaneously. We adopt the ansatz that fixed point $(\hat{x},\hat{y})=(x_{0},y_{0})+\mu(x_{1},y_{1})+\mu^{2}(x_{2},y_{2})+\cdots$. We find the zeroth order solution as $(x_{0},y_{0})=(1/2,1/2)$ for which $f_{1}(y_{0})=0$ and $g_{1}(x_{0})=0$. Proceeding further, up to second order in $\mu$ we perturbatively find
$\hat{x}=x_{0}-\mu\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}+\frac{\mu^{2}}{g_{1}^{\prime}(x_{0})}\left[\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}g_{2}^{\prime}(y_{0})-\frac{1}{2}\left(\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}\right)^{2}g_{1}^{\prime\prime}(x_{0})\right]+\mathcal{O}(\mu^{3}),$ (83a)
$\hat{y}=y_{0}-\mu\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}+\frac{\mu^{2}}{f_{1}^{\prime}(y_{0})}\left[\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}f_{2}^{\prime}(x_{0})-\frac{1}{2}\left(\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}\right)^{2}f_{1}^{\prime\prime}(y_{0})\right]+\mathcal{O}(\mu^{3}).$ (83b)
Here, prime on a function denotes derivative with respect to the argument of that very function. Rather explicitly, we have
$(\hat{x},\hat{y})=\left(\frac{1}{2}-\frac{\mu}{28}-\frac{377\mu^{2}}{13720},\frac{1}{2}+\frac{\mu}{20}-\frac{193\mu^{2}}{7000}\right).$ (84)
Furthermore, on linearizing about this fixed point, the Jacobian is of the form
$J=\begin{bmatrix}\mu\hat{x}(1-\hat{x})f_{2}^{\prime}(\hat{x})&\hat{x}(1-\hat{x})f_{1}^{\prime}(\hat{y})\\ \hat{y}(1-\hat{y})g_{1}^{\prime}(\hat{x})&\mu\hat{y}(1-\hat{y})g_{2}^{\prime}(\hat{y})\end{bmatrix},$ (85)
whose eigenvalues can be calculated, up to first order of $\mu$, to be $-\mu\pm({i}/{4})\sqrt{35+{26\mu}/{35}}$. Thus, the fixed point is a (locally asymptotically) stable spiral.
The neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ [i.e., Eq. (84)] is given as $(\mathbf{x},\mathbf{y})=((x,1-x),(y,1-y))=((\hat{{x}}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{{y}}+\delta_{y},1-\hat{y}-\delta_{y}))$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small but can have positive or negative sign, allowing $(\mathbf{x},\mathbf{y})$ to cover an entire neighbourhood around the point $(\hat{\mathbf{x}},\hat{\mathbf{y}})$. After some calculations, we can find that
$\mu{\sf A}[\mathbf{\hat{x}};\mathbf{x},\mathbf{x}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y},\mathbf{y}]-\mu{\sf A}[\mathbf{x};\mathbf{x},\mathbf{x}]-{\sf B}\left[\mathbf{x};\mathbf{y},\mathbf{y}\right]$ $=$ $-5\delta_{x}\delta_{y}+\mathcal{O}(\delta_{x}\delta_{y}\mu),$ (86a)
${\sf C}[\mathbf{\hat{y}};\mathbf{x},\mathbf{x}]+\mu{\sf D}[\mathbf{\hat{y}};\mathbf{y},\mathbf{y}]-{\sf C}\left[\mathbf{y};\mathbf{x},\mathbf{x}\right]-\mu{\sf D}[\mathbf{y};\mathbf{y},\mathbf{y}]$ $=$ $+7\delta_{y}\delta_{x}+\mathcal{O}(\delta_{y}\delta_{x}\mu).$ (86b)
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (86) is positive in the entire neighbourhood; thus, the criteria of or-definition is satisfied but it does not meet the criteria of ESS of and- and sum-definitions. Thus, the locally asymptotically stable fixed point given by Eq. (84) is 2ESS. But does it satisfy inequality (81)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
Since we forced to do numerical check, we no longer need to assume $\mu$ to be infinitesimal. In other words, we now consider Eq. (82) with positive real $\mu$ and find its interior fixed points numerically for various values of $\mu$. We also validate that all the fixed points are 2ESS as well. For each such fixed points, we check if inequality (81) is satisfied. To investigate that, we divide the entire $(x,y)$ space into a uniform grid of size $50\times 50$. We then determine the sign of minimum value of L.H.S. of inequality (81)—denoted as $\min_{\forall(x,y)}(-\dot{V_{r}})$—for different values of $\mu$ and $r$ in entire $(x,y)$ grid. Positive $\min_{\forall(x,y)}(-\dot{V_{r}})$ implies that inequality (81) is satisfied. As seen in Fig. 3, we find a (green) region (in $r$-$\mu$ space discretized as $5000\times 1000$ grid in numerics) bounded by black solid curves, where $\min_{\forall(x,y)}(-\dot{V_{r}})>0$; elsewhere, we find that $\min_{\forall(x,y)}(-\dot{V_{r}})<0$ (see the red region in Fig. 3). At each fixed point, we also do linear stability analysis and find that all the fixed points under question are locally asymptotically stable. In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 [inequality (81)]. We observe that if $\mu$ is very small, there is no such $r>0$ such that the inequality holds true for all $({x},{y})$; however, for higher values of $\mu$ such an $r$ exists.
Figure 3: Numerical demonstration that showing Proposition 1 is not always true, yet Conjecture 1 may be valid. We use payoff hypermatrices from Fig. 2 and multiply the parameter $\mu$ with matrices ${\sf A}$ and ${\sf D}$ to control the strength of intra-specific interactions. The figure illustrates two distinct regions: In the first region (green colour), inequality (81) in Proposition 1 is satisfied for all $(x,y)$ and in second region (red colour), it doesn’t. However, all the interior fixed points under consideration are checked to be locally asymptotically stable and 2ESS.
VII Conclusions
The primary objective of this study has been to critically revisit various definitions of ESS, specifically in the context of asymmetric (involving both inter and intra-specific interactions) and bimatrix games (with solely inter-specific interactions), and to unravel their suitability. Through our analyses, we have brought forward intricacies inherent in the definitions and have suggested their corresponding resolutions by examining them through the lens of dynamical stability. The insights, thus, gained resolve the prevailing ambiguities surrounding the ESS concept on asymmetric interaction scenarios.
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESS’s of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that an ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is non-strict NE. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed. While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
The most important take-home message, we would like to convey is that while strict NE is ESS under all three classes of definitions, other ESS’s need not simultaneously satisfy all the classes. Not only that, the latter ESS’s may not always be locally asymptotically stable under replicator equation. In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose. The asymptotic approach towards ESS state is also compatible with the principle of minimum relative entropy—a rather general principle of nature.
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Establishing pure 2ESS’s asymptotic stability in multiplayer games using techniques beyond linear stability analysis is also missing. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2023), opens up another avenue for investigation.
Acknowledgements.
Authors thank Ross Cressman for helpful remarks on the initial idea of this work. Arunava Patra thanks CSIR (India) for the financial support in the form of Senior Research Fellowship. Sagar Chakraborty acknowledges the support from SERB (DST, govt. of India) through project no. MTR/2021/000119.
Appendix A Comment on Ref. Cressman (1996)’s proof of Theorem 1
In Appendix A of Ref. Cressman (1996), Theorem 1 is seemingly proven in the context of asymmetric games which are defined to include bimatrix games. While the elaborate proof is best seen in the original reference, let us briefly point out at which point of argument the proof ceases to be valid for bimatrix games.
We begin by referring to a statement present in the proof: “When $\theta=0$, the point is$-((p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1}),(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ which is in quadrant IV. Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted as $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively. $A,\,B,\,C,\,D$ are payoff matrices ${\sf A,\,B,\,C,\,D}$. In this proof, one tries to show that Definition 3d implies inequality (46); for this purpose, one introduces $\theta$ as a parameter to explore in $\mathbb{R}^{2}$ space spanned by $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$ and $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ as the two coordinates. Idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (46) is satisfied.
A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed NE), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero. Let us see explicitly why this is so.
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero-sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
In conclusion, if one accepts the theorem—present in Ref. Cressman (1996)—exactly as worded, then one is prone to erroneously deducing that mixed 2ESS in bimatrix games is asymptotically stable. This oversight in Ref. Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
We now demonstrate that there cannot be a mixed ESS under Definition 2e in bihypermatrix games (${\sf A=D=0}$). While we focus on the 2-strategy case for simplicity, the arguments can be easily generalized for any number of strategies. We begin by noting that for mixed ESS state, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following condition holds [see Eq. (78)]:
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]~{}\text{and}~{}{\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (87)
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\delta}_{x},\mathbf{\hat{y}}+\bm{\delta}_{y})\equiv((\hat{x}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{y}+\delta_{y},1-\hat{y}-\delta_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\delta}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\delta}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (88)
To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood. To this end, we expand the terms in Definition 2e to arrive at,
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ $=$ $(\hat{x}-x)\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)+(\hat{y}-y)\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\delta_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\delta_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\delta_{x}\left(b_{1}\delta_{y}+b_{2}\delta_{y}^{2}+\cdots+b_{\gamma_{x}}\delta_{y}^{\gamma_{x}}\right)-\delta_{y}\left(c_{1}\delta_{x}+c_{2}\delta_{x}^{2}+\cdots+c_{\nu_{x}}\delta_{x}^{\nu_{y}}\right).$
To get to the final form Eq. (LABEL:eqn:bihyp_ESS_s), we have used Eq. (87) and Eq. (88). The coefficients $b_{1},b_{2},\cdots,b_{\gamma_{x}}$ depend on elements of ${\sf B}$ and $\hat{y}$; similarly, $c_{1},c_{2},\cdots,c_{\nu_{y}}$ depend on elements of ${\sf C}$ and $\hat{x}$.
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\delta_{x}\delta_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\delta_{x}$ and $\delta_{y}$. For similar reason, in case $-(b_{1}+c_{1})\delta_{x}\delta_{y}=0$, the next order term in $\delta_{x}$ and $\delta_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.